OpenStudy (anonymous):

So I have a question and I am not sure what it means... Compare P(A∩B) with P(A∪B), and explain what each probability means in the context of the situation and data you collected.

OpenStudy (anonymous):

@ganeshie8 the table is the same: Type krispy kreme dunkin donuts home 5 2 bakery 7 6

ganeshie8 (ganeshie8):

easy, define \(A\) and \(B\) first

ganeshie8 (ganeshie8):

lets say \(A\) : krispy kreme \(B\) : home

ganeshie8 (ganeshie8):

\(P (A \cap B)\) means, "Probability of both A and B occurring"

ganeshie8 (ganeshie8):

look at the table, how many like both "krispy kreme" and "home" ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

I am sorry @ganeshie8! I got busy with something else and took a lunch break. I am sorry! Thank you for replying and helping! When you get back on, can you continue to help me with this please?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

So the answer for P(A∩B), it would be 5/12?

ganeshie8 (ganeshie8):

it should be 5/20

ganeshie8 (ganeshie8):

you need to divide by total number of students

ganeshie8 (ganeshie8):

and, P(A∪B) represents "probability for either A or B or both"

OpenStudy (anonymous):

Oh okay so the answer for P(A∩B) is 1/4 when simplified. So then would P(A∪B) be 1/4 or 12/20 (or 3/5)? I get 12/20 because I added the home and the bakery under krispy kreme. Am I doing this correct?

OpenStudy (anonymous):

12/20 simplifies to 3/5, that is where 3/5 came from

ganeshie8 (ganeshie8):

denominator is right

ganeshie8 (ganeshie8):

how did u get 12 for numerator ?

OpenStudy (anonymous):

Okay :) I added 5 (home) + 7 (bakery) to get 12 total under the krispy kreme row.

OpenStudy (anonymous):

I mean under the krispy kreme column

ganeshie8 (ganeshie8):

why did u leave "home AND dunken donuts" ?

ganeshie8 (ganeshie8):

u want either "home" or "kirspy kreme"

ganeshie8 (ganeshie8):

so A U B includes everyone in home, and everyone in kirspy kreme

OpenStudy (anonymous):

So how do we solve it?

ganeshie8 (ganeshie8):

everyone in home : 5 + 2 = 7 everyone in krispy kreme : 5 + 7 = 12 add them : 19 however we have added 5 times, so subtract 5 : 19 - 5 = 14

ganeshie8 (ganeshie8):

so P(A U B) = 14/20

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

Oh okay. So simplified it would be 7/10?

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

okay! Thank you so much! I get it now! :)

OpenStudy (anonymous):

Does this sound good? A : krispy kreme B : home P(A∩B) means Probability of both A and B occurring In the home column under Krispy Kreme, we have 5. This will be the numerator since we are finding the probability of people who eat Krispy Kreme at home as opposed to the probability of people eating Krispy Kreme at the Bakery, eating Dunkin Donuts at Home, and eating dunkin donuts at the bakery. So the denominator is going to be 20 since 5 + 7 = 12, 12 + 2 = 14, 14 + 16 = 20 P(A∪B) represents "probability for either A or B or both. So what we do is find the sum of the home row. 5 + 2 = 7 Then we find the sum of the Krispy Kreme column. 5 + 7 = 12 Then we add the sums together. 7 + 12 = 19 Then, since we added 5 to both, we have to subtract 5. 19 - 5 = 14 Because the sum of all of the people total in the table is 20, we have to put 14 over 20. As you can see, we can simplify this to be 7/10.

ganeshie8 (ganeshie8):

looks good !!

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

Thank you!!

ganeshie8 (ganeshie8):

u wlc !